• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Techs

    Log entry time 16:02:47 on August10,2006

    Entry number 178884

    keyword=Left Q2 and bogie motors

    Nitrogen meter on Left Q2 was ok. The control valve was stuck open. Mark set the return valve to try and control the flow. The bogie motors on the Left Arm seem to be ok. It looks like the wheels were slipping. With a little care one can get the spectrometer to where it needs to go. 700 RPM seems to be a good speed for the motors. Don't set them below 300 RPM.